# HLStrans: Dataset for LLM-Driven C-to-HLS Hardware Code Synthesis

### v1

@@ 4.2.1 Monte Carlo tree search (MCTS)
Agent G: We propose to use an LLM as the agent that leverages its pretrained knowledge of hardware design and in-context learning abilities.
Action A: At each time step $t$, the agent selects an action $a_{t}$, which corresponds to a prompt or transformation applied to the current HLS code. We define two complementary action types: RAG-based actions retrieve optimization policies directly from our pre-built table and accompanying code examples shown in Figure 4, leveraging retrieval-augmented generation to surface proven transformations rapidly and reliably. Reasoning-based actions with ReAct prompt yao2023react , in contrast, analyze compiler warnings such as pipeline-interval breaches or loop-unroll violations and apply targeted code reforms by interpreting warning semantics within the current code context.
(a) Loop tiling Code examples by RAG
(b) LLM reasoning about environment warning and tool hint
Figure 4: Actions design of MCTS
Reward R: Rule-based reward shaping has proven effective in guiding agent behavior in previous work guo2025deepseek . In our framework, the reward function $R(s_{t},a_{t})$ is computed by applying rule-based scoring to verification results and feedback provided by the HLS tool. A penalty of $-2$ is applied when the verification fails, and $-1$ if synthesis fails or the design exceeds resource constraints. A neutral reward of 00 is given when the transformed design performs worse than the original, while a reward of $1$ is granted when it performs better. If the design not only surpasses the original but also meets timing constraints, a higher reward of $2$ is assigned.
In our cases, the MCTS begins at the initial state $S_{0}$ (the root node), which is the naive HLS code. From a state $S_{t}$, the agent applies an optimization policy $\pi$, i.e., an action $a_{t}\in\mathcal{A}$, transitioning to the subsequent state $S_{t+1}$. MCTS consists of four key phases: selection, expansion, simulation, and backpropagation.
Selection: We employ the upper confidence bounds for trees (UCT) gelly2006exploration algorithm to choose nodes.
Expansion: From the current state $s_{t}$, generate one or more child nodes to explore untried actions, formalized as $a_{t+1}=\mathrm{llm}(s_{t})$.
@@ 4.2.2 Design Space Exploration
The tool implements an automated HLS design-space explorer that uses a genetic-algorithm optimizer to discover effective directive combinations, specifically loop pipelining, loop unrolling, and array partitioning, that maximize performance and resource utilization. By encoding each candidate solution as a set of directive parameters, evolving them through mutation and crossover, and evaluating them via HLS synthesis feedback, the tool generates high-quality configurations without manual tuning.
Overall, we leverage DeepSeek-R1 to generate high-quality synthetic code examples, the AMD Vitis HLS EDA tool, and DSE tools to validate, annotate, and collect performance/resource metrics within our framework, yielding an effective, end-to-end HLS code dataset. The detailed information on dataset generation is in Appendix A.2.
@@ 5.1 Prompting Methods
Zero-shot Prompting: We craft concise, HLS-specific prompts that instruct the model to perform code optimization or transformation from its pretrained knowledge, without any additional fine-tuning or example demonstrations liu2021makes wei2021finetuned .
Chain-of-Thought Prompting: Building on the chain-of-thought approach of Wei et al. wei2022chain , our prompts first guide the model through a transformation reasoning phase before asking it to emit the refined code.
Retrieval-Based Prompting: Recent studies shrivastava2023repository shypula2023learning have shown that retrieval-based techniques can substantially boost code generation quality in large language models. In our approach, we first encode each program using CodeBERT zhou2023codebertscore to produce rich, semantically informed embeddings. We then index these vectors with FAISS johnson2019billion (Facebook AI Similarity Search) and perform a K-nearest-neighbors lookup to retrieve the top K most similar code snippets from our training corpus. Finally, we supply these retrieved examples alongside the original code as additional context to the LLM, guiding it to produce more accurate and effective edits. In our experiments, we set K to 1. The detailed prompt information is in Appendix A.3.
@@ 5.2 Experiment setting
Fine tuning setting. Training was conducted using 2 NVIDIA H100 GPUs, each with 80 GB of memory. The computing environment was configured with CUDA 12.2 and cuDNN 9.1 to ensure optimal deep learning performance.
Optimization Conditioned Fine-tuning: Programs that solve the same task can vary widely in how well they perform because of many different kinds of directives and code transformations. The optimized code has a jumble of small or large speedups, without any signal about which gains are most significant. To give the model clearer guidance, we borrow ideas from shypula2023learning and attach a “performance tag” and "resource tag" to each solution during training. Each tag reflects how close that program comes to the best attainable performance with resources minimized, using a discrete scale from 1 to 10, respectively. Concretely, for every benchmark problem, we rank its optimized solutions by latency and resource, labeling the top 10% in performance as a “10/10” performance tag and the lowest 10% in resource usage as a “10/10” resource tag. By examining these tags, the model learns which code patterns tend to yield the highest performance. At inference time, we prompt the model to generate the best-performing codes with the least resources. The detailed fine-tuning information is in Appendix A.4.
Models. We evaluate closed-source models, GPT-4o and deepseek-R1. We also used pretrained checkpoints of open-source Qwen models yang2024qwen2 : Qwen2.5-Coder-32B-Instruct, Qwen2.5-Coder-7B-Instruct, and Qwen2.5-Coder-3B-Instruct obtained via HuggingFace wolf2020transformers .
Metrics. To evaluate the effects of HLS code generations, we use the following metrics:
  • Functional Accuracy (%Function Correct): The share of test programs that preserve the original functionality testbench.
  • Synthesis Accuracy (%Synthesis Correct): The share of programs that can be synthesized.
  • Reduce Latency (Acceleration): The ratio of original latency to new latency, $accel=o/n$, where $o$ is the old latency and $n$ is the new latency from reports of HLS synthesis respectively. We report the minimum, average, and maximum acceleration only for the functionally and syntactically correct cases.
  • Optimization Rate (%OPT): The percentage of test programs that reduce latency (Acceleration > 1) while remaining functionally correct.
FPGA Flow Setting: We transform the given sequential C++ code of these benchmarks into optimized HLS kernel code. We adopt the Vitis HLS design flow, using the Xilinx Alveo U55C data center accelerator card. The target frequency is 200 MHz.
We selected 17 applications as test cases both unsynthesizable designs requiring repair and synthesizable designs in need of optimization and used the remaining 120 applications for training (90%) and validation (10%). Previous works li2022competition show that generating multiple program candidates per input and selecting the optimal one improves code synthesis performance. We generate $k$ program variants for each input, then select the fastest one that successfully passes all test cases; we refer to this sampling-and-selection strategy as $Best@k$.
@@ 5.3.1 Results of Pretrained Models
Table 3 and 4 report Best@1 and Best@5 accuracy for different prompts. Overall, deepseek-R1 achieves the highest performance, and every model benefits from chain-of-thought prompts compared to zero-shot. However, we observe that some LLM-optimized codes actually degrade performance. This can happen for two reasons: First, the restructuring performed by the LLM can introduce new loop dependencies, increasing latency; Second, the pragmas inserted by LLMs may be less effective than the default optimizations inferred by the HLS compiler.
Table 3: Best@1 Results for various methods and models.
Acceleration
Method Model Opt Min Avg Max Function correct Synthesis correct
Zero-shot Deepseek-R1 17.6% 0.3× 2.8× 24.6× 47.1% 29.4%
GPT-4o 17.6% 0.5× 1.6× 10.9× 35.3% 29.4%
Qwen coder 32B 5.9% 0.8× 1.3× 1.5× 41.2% 11.8%
Qwen coder 7B 0% 0.66× 0.9× 1.0× 38.8% 11.8%
Qwen coder 3B 0% 0.75× 0.8× 1.0× 11.8% 11.8%
COT Deepseek-R1 17.6% 0.9× 4.3× 41.0× 52.9% 29.4%
GPT-4o 29.4% 0.5× 1.3× 4.0× 52.9% 35.3%
Qwen coder 32B 11.8% 0.1× 2.4× 24.6× 58.8% 17.6%
Qwen coder 7B 17.6% 0.04× 1.2× 2.0× 52.9% 35.3%
Qwen coder 3B 5.9% 0.8× 1.1× 1.8× 47.1% 29.4%
Retrieval Prompt Deepseek-R1 41.2% 0.9× 7.2× 57.0× 47.1% 47.1%
GPT-4o 41.2% 0.9× 4.7× 24.6× 58.8% 47.1%
Qwen coder 32B 29.4% 0.7× 4.9× 24.6× 47.1% 35.3%
Qwen coder 7B 17.6% 0.3× 2.1× 5.1× 52.9% 35.3%
Qwen coder 3B 11.7% 0.7× 1.0× 3.4× 29.4% 11.8%
@@
Retrieval-augmented prompting consistently outperforms purely generative methods, achieving the highest synthesis success rates without degrading functional correctness. This demonstrates that our dataset by providing a rich cache of validated pragmas and code transformation examples serves as an indispensable “best practices” repository, steering LLMs toward hardware-friendly idioms and dramatically reducing synthesis failures.
Table 4: Best@5 Results for various methods and models.
Acceleration
Method Model Opt Min Avg Max Function correct Synthesis correct
Zero-shot Deepseek-R1 23.5% 0.8× 4.3× 23.5× 47.1% 29.4%
GPT-4o 23.5% 0.7× 2.4× 23.5× 41.1% 29.4%
Qwen coder 32B 6.9% 0.9× 2.1× 20.4× 64.7% 35.3%
Qwen coder 7B 0% 0.05× 0.5× 1.0× 47.1% 17.6%
Qwen coder 3B 5.9% 0.59× 1.4× 8.0× 52.9% 23.5%
COT Deepseek-R1 23.5% 0.5× 8.2× 53.2× 52.9% 41.2%
GPT-4o 29.4% 0.7× 3.1× 17.1× 52.9% 35.3%
Qwen coder 32B 17.6% 0.5× 3.5× 20.4× 52.9% 35.3%
Qwen coder 7B 17.6% 0.1× 0.7× 1.0× 47.1% 23.5%
Qwen coder 3B 11.8% 0.6× 1.5× 2.3× 47.1% 35.3%
Retrieval Prompt Deepseek-R1 47.2% 0.9× 9.4× 83.2× 64.7% 58.8%
GPT-4o 41.2% 0.6× 5.9× 37.3× 58.8% 47.1%
Qwen coder 32B 35.3% 1.2× 5.1× 24.6× 47.1% 35.3%
Qwen coder 7B 11.8% 0.5× 2.7× 5.1× 52.9% 11.8%
Qwen coder 3B 17.6% 0.6× 1.5× 3.2× 35.3% 29.4%
5.3.2 Results of Fine-Tuning Models
Both the 3B and 7B fine-tuned models show consistent gains in optimization quality, latency reduction, and synthesis success rate in Table 5. They generate HLS code that not only executes faster but also synthesizes more reliably, even though the function-correct rate has slightly dropped. These results demonstrate that training on our curated dataset significantly boosts an LLM’s ability to produce correct, high-performance HLS implementations directly from C sources.
Table 5: Fine-Tuning: Best@1 vs. Best@5 Results for Various Models
Acceleration
Setting Model Opt Min Avg Max Function correct Synthesis correct
Best@1 Qwen coder 3B 11.8% 0.1× 2.1× 7.7× 35.3% 35.3%
Best@1 Qwen coder 7B 17.6% 0.4× 2.8× 17.2× 41.1% 29.4%
Best@5 Qwen coder 3B 23.5% 0.6× 2.9× 8.9× 41.1% 35.3%
Best@5 Qwen coder 7B 29.4% 0.8× 3.1× 21.4× 41.1% 35.3%
@@ 6 Conclusion
We introduce a novel dataset that transforms C kernels into richly annotated HLS implementations, empowering LLMs to learn hardware-aware optimizations such as loop pipelining, unrolling, and memory buffering. Our experiments demonstrate that retrieval and fine-tuning on this dataset significantly boosts both latency reduction and synthesis success rates, proving its effectiveness in accelerating and automating electronic design flows. By releasing both the dataset and training scripts, we aim to catalyze further exploration at the intersection of LLMs and hardware design.
7 Future work
We evaluate HLS designs based on synthesis reports, which yield resource and timing estimates rather than true hardware performance. In future work, we will measure and compare real execution time across different models on physical FPGA platforms.
Unlike large software-optimization datasets, publicly available HLS benchmarks remain scarce. To address this gap, we will explore more effective methods for generating HLS-specific datasets and continue to grow our collection of real-world examples. Our current dataset targets AMD FPGAs; we plan to extend this work by generating analogous datasets for Intel FPGAs and Catapult HLS.
Appendix A Appendices

### v3

@@ 3.2.1 Monte Carlo tree search (MCTS)
We formulate HLS optimization as an MCTS problem. The environment is the Vitis HLS toolchain, which provides synthesis, resource, and performance feedback. The agent is an LLM that applies code transformations. Actions include (i) RAG-based retrieval of known optimization policies and (ii) ReAct-based reasoning over compiler warnings. The state is the current HLS code, and the reward follows rule-based shaping: 2-2 for verification failure, 1-1 for synthesis/resource failure, 0 if worse, $1$ if improved, and $2$ if improved with timing met. In our cases, the MCTS agent begins at the initial state $S_{0}$ (the root node), which is the naive HLS code. From a state $S_{t}$, the agent applies a optimization policy $\pi$, i.e., an action $a_{t}\in\mathcal{A}$, transitioning to the subsequent state $S_{t+1}$. This new state optimizes the existing code sequence by applying the new optimizations. Upon reaching a terminal state $S_{T}$, the agent receives a deferred reward $R(S_{T})$. $N(S_{t})$, the total number of times $S_{t}$ has been visited.
Selection: We employ the upper confidence bounds for trees (UCT) algorithm (Gelly & Wang, 2006) to choose nodes. The UCT formula includes the average reward for the current state, which encourages the path that can bring high reward, while $U$ term measures the associated uncertainty, which encourages the exploration of new paths. This approach effectively balances the trade-off between exploration and exploitation.
Expansion, simulation and backpropagation: Expansion is to explore the unchosen action. We leverage LLM to determine the next action from the unexplored. The decision process is driven by program analysis in conjunction with the history of adopted optimizations, enabling LLM to accurately assess and select the most promising action. After the analysis of LLM for state $s_{t}$ at time steps $t$, the next action $a_{t+1}$ will decided by $a_{t+1}=llm{(s_{t})}$. Simulation employs the agent to apply transformations and evaluates them via HLS synthesis measuring estimated latency and resource usage to compute the reward $R(s_{t},a_{t})$; During backpropagation, these rewards are propagated up the search tree to update node values, refining the agent’s estimates and guiding future action selection. Once we no longer observe significant improvements, the search process is halted, and the best-performing rewritten design is selected. The detailed description of MCTS framework is described in Appendix A.2.
@@ 3.2.2 Design Space Exploration
The tool implements an automated HLS design-space explorer that uses a genetic-algorithm optimizer to discover effective directive combinations, specifically loop pipelining, loop unrolling, and array partitioning, that maximize performance and resource utilization. To traverse the solution space, we utilize the NSGAII algorithm (Deb et al., 2002) implemented in PyMOO library (Blank & Deb, 2020), known for its ability to bypass local optimal and quickly converge to efficient solutions. The detail DSE implementation is introduced in Appendix A.2.2.
3.3 Dataset Selections
After generating multiple dataset candidates, we select the efficient samples. If the input codes can not be synthesized, we choose the candidates which can be synthesized. If the input codes can be synthesized, we choose the candidates whose latency is less than input codes. To give the model clearer guidance, we borrow ideas from Shypula et al. (2023) and attach a “performance tag” and ”resource tag” to each solution during training. Each tag reflects how close that program comes to the best attainable performance with resources minimized, using a scale from 0 to 10, respectively.
3.4 HLStrans Statics
Overall, we leverage DeepSeek-R1 to generate high-quality synthetic code examples, the AMD Vitis HLS EDA tool, and DSE tools to validate, annotate, and collect performance/resource metrics within our framework, yielding an effective, end-to-end HLS code dataset. Our dataset has the following merits:
Diverse Application Coverage. Table 1 shows that HLStrans provides the largest number of HLS kernels and the longest average lines of code, incorporating commonly used HLS benchmarks as well as real-world examples. Our curated corpus spans diverse application domains and covers all six transformation categories listed in Appendix A.1. Figure 4(a) visualizes the program distribution across these five tasks, and the source kernels themselves fall into seven distinct application categories. This rich, well-balanced dataset offers broad coverage of real-world HLS patterns required to train and evaluate LLMs’ hardware-synthesis capabilities.
Diverse types of transformations. To evaluate the LLM’s ability to assess different C/C++ to HLS transformations, every transformation shown in Figure 1 must be supported. Each dataset sample may correspond to one or more types of transformations. Figure 4(b) illustrates how the number of samples for each transformation increases after data augmentation.
Figure 4: (a) Program source distribution. (b) Percentage of different transformations. (c) Speedup percentiles across dataset.
High quality of dataset samples. To evaluate the quality of the dataset, we measured the execution-cycle ratio between the original and target codes using reports from Vitis HLS. The speedup is ratio between the latency of the original design and the generated design, as reported by the synthesis tool. We then computed the percentile distribution of these speedup values across all pairs. As shown in the Figure 4(c), 100% of the pairs, the target code is $\geq 1.5\times$ faster, and for 25% of the pairs, it achieves a speedup of $\geq 50.3\times$. Different samples are annotated with performance and resource usage tags, allowing the LLM to understand the detailed effects of C-to-HLS transformations. This enables the LLM to generate code that achieves higher performance while consuming fewer resources. The detailed information on dataset generation is in Appendix A.2. The datasets are released under the MG0‑2.0 Non‑Commercial (NC) license  (Duan et al., 2024). [^1] https://www.modelgo.li/.These licenses permit both academic and commercial reuse provided that attribution is given. The dataset release also includes provenance metadata and third-party license notices.
@@ 4 Experimental Evaluation of LLMs on Our Dataset
To evaluate LLM performance on our dataset and assess the dataset’s impact on model capability, we explore different prompting strategies and fine-tune smaller models using supervised fine-tuning (SFT) (Ouyang et al., 2022).
@@ 4.1 Prompting Methods
Zero-shot Prompting: We craft concise, HLS-specific prompts that instruct the model to perform code optimization or transformation from its pretrained knowledge, without any additional fine-tuning or example demonstrations (Liu et al., 2021) (Wei et al., 2021). Chain-of-Thought Prompting: Building on the chain-of-thought approach of Wei et al. (2022), our prompts first guide the model through a transformation reasoning phase before asking it to emit the refined code. Retrieval-Based Prompting: Recent studies (Shrivastava et al., 2023) (Shypula et al., 2023) have shown that retrieval-based techniques can substantially boost code generation quality in large language models. In our approach, we first encode each program using CodeBERT (Zhou et al., 2023) to produce rich, semantically informed embeddings. We then index these vectors with FAISS (Johnson et al., 2019) (Facebook AI Similarity Search) and perform a K-nearest-neighbors lookup to retrieve the top K most similar code snippets from our training corpus. Finally, we supply these retrieved examples alongside the original code as additional context to the LLM, guiding it to produce more accurate and effective edits. In our experiments, we set K to 1. The detailed prompt information is in Appendix A.3.
@@ 4.2 Experiment setting
To evaluate LLMs on our dataset, we have the following evaluation setting. Task setup: Given a C/C++ kernel, the model must generate an optimized HLS implementation. Success requires not only functional correctness but also synthesizability under FPGA toolchains. Models: We benchmark both closed-source (GPT-5 (Wang et al., 2025), DeepSeek-R1 (Chua & Evans, 2025), Grok 4 (xAI, 2025), Gemini 2.5 Pro(Gemini Team, 2025)) and open-source (Qwen 2.5 Coder (Hui et al., 2024)) models, under different prompting strategies (zero-shot, chain-of-thought, retrieval) and fine-tuning. Dataset split: Following standard machine-learning protocol, we reserve 270 applications for training and validation and hold out 39 applications for evaluation. The held-out set includes both unsynthesizable designs that require repair and synthesizable designs that require optimization. Crucially, these 39 held-out applications were excluded from the LLM-based data-augmentation pipeline to prevent any risk of data leakage into the evaluation. Infrastructure: All synthesis is conducted with the Xilinx Vitis HLS toolchain targeting a datacenter FPGA (Alveo U55C). Training was conducted using 2 NVIDIA H100 GPUs, each with 80 GB of memory. The computing environment was configured with CUDA 12.2 and cuDNN 9.1 to ensure optimal deep learning performance. Metrics: Unlike conventional code generation benchmarks that stop at functional correctness, the C-to-HLS task requires models to satisfy both software and hardware constraints. We therefore report four complementary metrics:
  • Functional Accuracy: The share of test programs that preserve the original functionality testbench.
  • Synthesis Accuracy: Percentage of programs that compile successfully into FPGA-ready hardware.
  • Speedup (Latency reduction): Ratio between the latency of the original design and the generated design, as reported by the synthesis tool.
  • Optimization Rate (%OPT): Fraction of generated programs that both pass correctness checks and achieve speedup $1\times$.
Previous works (Li et al., 2022) show that generating multiple program candidates per input and selecting the optimal one improves code synthesis performance. We generate $k$ program variants for each input, then select the fastest one that successfully passes all test cases; we refer to this sampling-and-selection strategy as $Best@k$.
@@ 4.3 Experiment Results
4.3.1 Evaluation of Data Augmentation Framework.
The MCTS component of our framework can produce variable iteration lengths. To quantify this behavior, we evaluated both runtime and achieved speedup on the PolyBench suite (Pouchet & Yuki, 2012) while sweeping the number of rollouts. Figure 5(a) reports these results and indicates that 32 rollouts is the “sweet spot”; consequently, we set the rollout count to 32 for subsequent experiments. We also compared our framework against state-of-the-art approaches on the Rodinia benchmarks by measuring the runtime of the optimized programs on real FPGA cards. Figure 5(b) shows the runtime comparison for five Rodinia benchmarks (Che et al., 2009). With the same base model, our framework attains more than $5\times$ average speedups than Xiong et al. (2024). We additionally observed that DeepSeek-R1 produces even better results; therefore, we selected DeepSeek-R1 as the generator for new data used in later experiments. These findings motivated both our choice of rollout parameter and our selection of the data-generation model. The detailed experiments comparison results are in Appendix A.2.
(a) The effect of MCTS rollout setting
(b) Comparison with sota framework
Figure 5: Evaluation of our dataset augmentation framework (a) MCTS rollout setting. (b) Rollout comparison.
4.3.2 Results of Fine-Tuning Models
Both the Qwen2.5 Coder 3B and 7B fine-tuned models show consistent gains in optimization quality, latency reduction, and synthesis success rate in Table 2. They generate HLS code that not only executes faster but also synthesizes more reliably, even though the function-correct rate has slightly dropped. These results demonstrate that training on our curated dataset significantly boosts an LLM’s ability to produce correct, high-performance HLS implementations directly from C sources.
Table 2: Fine‐Tuning results comparison. Transformation: the T1–T5 transformation in Figure 1 applied to examples that are functionally and synthesis correct.
Speedup Transformation
Method Model Opt Min Avg Max T1 T2 T3 T4 T5
Functional
Accuracy
Synthesis
Accuracy
[…]
SFT
[…]
Efficient HLS kernels require a mix of C to HLS transformations. We measure how our dataset improves LLM C to HLS optimization: Table 2 demonstrates that fine-tuning on our corpus raises success rates across transformation types.
@@ 4.3.3 Results of Pretrained Models
Table 3 reports the Best@1 and Best@5 accuracies for different prompts and models. To evaluate the utility of our dataset, we constructed retrieval databases from the HLSdataset (Wei et al., 2023) and from our training data, and applied retrieval-based prompting using these two databases to measure its effect. Overall, incorporating our dataset into retrieval improved pretrained models performance compared with other prompt methods.
Table 3: Best@1 and Best@5 results for various methods and models.
Best@1 Best@5
Speedup Speedup
Method Model Opt Min Avg Max
Functional
Accuracy
Synthesis
Accuracy
Opt Min Avg Max
Functional
Accuracy
Synthesis
Accuracy
[…]
Retrieval Prompt
[…]
Retrieval Prompt
[…]
4.4 Results Analysis
@@
Observation 1. Retrieval-augmented generation and finetuning on HLStrans can improve model’s performance on C to HLS task. This demonstrates that our dataset by providing a rich cache of validated pragmas and code transformation examples serves as an indispensable “best practices” repository, steering LLMs toward hardware-friendly idioms and dramatically reducing synthesis failures.
Observation 2. Sampling diversity substantially boosts results. Allowing up to five candidate generations per input (Best@5) improves both synthesis success rate and achieved acceleration. For example, GPT-5 (retrieval prompt with HLStrans) raises synthesis accuracy from 53.8% to 71.8% under Best@5, underscoring the benefit of n-best generation.
Observation 3. LLM optimization may harm HLS code performance. We observe that some LLM-optimized kernels actually degrade performance (speedup $1\times$). This can happen for two reasons: First, the restructuring performed by the LLM can introduce new loop dependencies, increasing latency; Second, the pragmas inserted by LLMs may be less effective than the default optimizations inferred by the HLS compiler. Therefore, it is necessary to set up dataset to guide LLM’s proper optimizations.
Observation 4. Trade-off between pass rate and optimizations. Applying retrieved, optimized code examples increases performance but reduces both functional and synthesis accuracy. For example, Deepseek-R1 (retrieval prompt with HLSdataset) increase speedup but decrease functional accuracy from 43.6% to 33.3% compared with zero-shot prompt. This highlights a trade-off between aggressive optimization and correctness when the LLM’s capability is unchanged.
Observation 5. LLMs perform differently across transformations. As shown in Figure 1, pretrained models more easily apply T2 and T5. Fine-tuning on HLStrans improves the success rates of all the transformations, as reported in Table 2.
@@ 5 Conclusion
We introduce a novel dataset that transforms C or C++ kernels into richly annotated HLS implementations, empowering LLMs to learn hardware‑aware optimizations such as loop pipelining, unrolling, and memory buffering. Our experiments demonstrate that retrieval and fine‑tuning on this dataset significantly boosts both latency reduction and synthesis success rates, proving its effectiveness in accelerating and automating electronic design flows. By releasing the dataset and training scripts, we aim to catalyze further exploration at the intersection of LLMs and hardware design.
Ethics Statement
We release a dataset that converts C/C++ kernels into richly annotated HLS implementations, together with training scripts, to accelerate LLM-driven hardware optimizations. While retrieval and fine-tuning improve latency and synthesis success, automated optimizations can produce incorrect or biased transformations; therefore the dataset and models are for research-only use and not intended for safety-critical deployment. Users should apply human review, evaluate functional correctness and synthesis safety alongside performance gains, and publish datasheets/model cards to promote transparency. Continued work on verification, robustness, and responsible reporting of failure cases is strongly encouraged.
Reproducibility Statement
We are committed to ensuring the reproducibility of our findings. All datasets, code, and experimental scripts are publicly available at https://anonymous.4open.science/r/HLStrans-B578/.
LLM Usage Declaration
We used Gemini 2.5 Pro [^2] https://deepmind.google/models/gemini/pro/ to polish grammar and phrasing during the writing process. No part of the analysis, experimental design, or results was generated by a large language model.
Appendix A Appendix